OpenStudy (anonymous):

Find the sum of the infinite geometric series, if it exists. 4 - 1 +1/4 - 1/16 + . . . A. - 1 B. 3 C. D. does not exist

OpenStudy (mathmale):

Yes, this is a geometric series. What is the common ratio, r? Is the absolute value of this r less than 1? What is the formula for the sum of a geom. series when the first term and the common ratio r are known?

OpenStudy (anonymous):

HI :DD

OpenStudy (anonymous):

\[ \Large \sum_{i=0}^k a r^i= \frac{a \left(1- r^{k+1}\right)}{1-r} \]

OpenStudy (anonymous):

If |r| < 1 \[ \Large \sum_{i=0}^\infty a r^i= \frac{a }{1-r} \]

OpenStudy (anonymous):

Your a is 4 and your \( \Large r = -\frac 1 4 \)

OpenStudy (anonymous):

Replace in the above formula and you are done.

OpenStudy (anonymous):

i cant understand the formula probably becausse how its written

OpenStudy (anonymous):

Read this http://mathworld.wolfram.com/GeometricSeries.html

OpenStudy (anonymous):

What is the choice C

OpenStudy (anonymous):

16/5 is c

OpenStudy (anonymous):

@eliassaab is it c?

OpenStudy (anonymous):

Yes it is 16/5

OpenStudy (anonymous):

thank you , could you help me with another please?

OpenStudy (anonymous):

Post it as a new problem and tag me in it

OpenStudy (raden):

just an alternative : let A = 4 - 1 +1/4 - 1/16 + . . . A = -4(-1 + 1/4 - 1/16 + . . .) see if A = 4 - 1 +1/4 - 1/16 + . . . then A - 4 = - 1 +1/4 - 1/16 + . . . so, A = -4 (A - 4) A = -4A + 16 group them, get 5A = 16 A = 16/5 sometimes, algebraic is enjoy :)

OpenStudy (anonymous):

thank you @RadEn

OpenStudy (raden):

welcome :)